OpenStudy (anonymous):

Solve by using the square root method, express answer in simplest radical form: 3x squared+96=0

OpenStudy (campbell_st):

divide every term by 3... as the 1st step, what do you get..?

OpenStudy (anonymous):

should i subtract 96 first though?

OpenStudy (anonymous):

@campbell_st ^^^

OpenStudy (anonymous):

i got for my answer 4i radical 2

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

the answer is correct... x^2 + 32 = 0 x^2 = -32 \[x = \pm 4i \sqrt{2}\]

OpenStudy (anonymous):

thanks :)